OpenStudy (anonymous):

A square based container has side 12cm and height 18 m. what is its lateral surface area?

OpenStudy (anonymous):

the area is 576?

OpenStudy (lgbasallote):

\[\huge LSA = 4sh\] where: s = side h = height i don't think that's what you did

OpenStudy (anonymous):

ahh. lateral surface area is differ from the normal surface area?

OpenStudy (aripotta):

yea, lateral area like, wraps around the figure. it's the surface area minus the bases

OpenStudy (anonymous):

what about the lateral surface area of rectangle?

OpenStudy (aripotta):

2d figures can't have lateral area

OpenStudy (aripotta):

there's area for 2d figures and lateral and surface area for 3d figures

OpenStudy (anonymous):

a rectangular prism is 5 m long, 3 m wide and 6 m high. what is its latera surdace area?

OpenStudy (anonymous):

126?

OpenStudy (aripotta):

SA rectangular prism = 2ab + 2bc + 2ac 2(5x3) + 2(3x6) + 2(5x6) = 30 + 36 + 60 = 126 m^2

OpenStudy (aripotta):

oh wait

OpenStudy (aripotta):

lateral area?

OpenStudy (anonymous):

lateral surface area

OpenStudy (aripotta):

LA = 2h(w+l) 12(3+5) = 12(8) = 96 m^2

OpenStudy (anonymous):

yehey!